The roots of x^2+kx+9=0 are real and equal. What are the possible values of k???
Answer by monika_p(71)

The roots of x^2+kx+9=0 are real and equal. What are the possible values of k???
If the roots are equal(double root) it means that discriminant of quadratic equation b^2-4ac=0
general form of quadratic equation is ax^2+bx+c=0
in this case a=1 b=k and c=9
b^2-4ac=0 then:
k^2-36=0
(k-6)(k+6)=0
k=6 or k=-6
For k=6 or k= -6 given equation has real and equal roots